Tuition costs at various colleges vary from $10,000 to $30,000 annually. These tuitions are expected to increase over time. If the annual rate of increase is 2 percent, what will be the new range in tuition costs in twelve years? Use Appendix A to answer the questions. Round your answers to the nearest dollar.
The range at 2 percent is $   - $   .

If the rate doubles from 2 to 4 percent, what will be the range in tuition costs after twelve years? Use Appendix A to answer the questions. Round your answers to the nearest dollar.
The range at 4 percent is $   - $   .
Answer #1
FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
a) FV = PV (1 + r )n
FV = 10000*(1+2%)^12
FV = 12682.42
FV = PV (1 + r )n
FV = 30000*(1+2%)^12
FV = 38047.25
Range = 12682.42 to 38047.25
b) FV = PV (1 + r )n
FV = 10000*(1+4%)^12
FV = 16010.32
FV = PV (1 + r )n
FV = 30000*(1+4%)^12
FV = 48030.97
Range = 16010.32 to 48030.97